A rational consumer spends income on two goods :bread and milk.when the consumer spends the last unit of income on bread total utility increases from 120 to 140 Utils.when the consumer spends the last unit of income of milk,total utility of milk increases from 50 to 60 utils.given the price of bread pb=4 dollar /unit ,find the price of milk,pm at the equilibrium point.
1
Expert's answer
2020-06-26T11:06:24-0400

"\frac{MUb}{Pb}" "=\frac{MUm}{Pm}"

"140-120=20" Utils of bread.

"60-50=10" Utils of milk

Price of bread "=" "4" dollars.

"\frac{20}{4}=\frac{10}{Pm}"

"Pm20=40"

"\text{The price of milk at equilibrium is 2 dollars}"